# From Robotics to Sepsis Treatment: Offline RL via Geometric Pessimism

### v1

@@ Abstract
Offline Reinforcement Learning (RL) promises the recovery of optimal policies from static datasets, yet it remains susceptible to the overestimation of out-of-distribution (OOD) actions, particularly in fractured and sparse data manifolds. Current solutions necessitates a trade off between computational efficiency and performance. Methods like CQL offers rigorous conservatism but require tremendous compute power while efficient expectile-based methods like IQL often fail to correct OOD errors on pathological datasets, collapsing to Behavioural Cloning. In this work, we propose Geometric Pessimism, a modular, compute-efficient framework that augments standard IQL with density-based penalty derived from $k$-nearest-neighbour distances in the state-action embedding space. By pre-computing the penalties applied to each state-action pair our method injects OOD conservatism via reward shaping with a $\mathcal{O}(1)$ training overhead.
Evaluated on the D4Rl MuJoCo benchmark, our method, Geo-IQL outperforms standard IQL on sensitive and unstable medium-replay tasks by over 18 points, while reducing inter-seed variance by 4 $\times$. Furthermore, Geo-IQL does not degrade performance on stable manifolds. Crucially, we validate our algorithm on the MIMIC-III Sepsis critical care dataset. While standard IQL collapses to behaviour cloning, Geo-IQL demonstrates active policy improvement. Maintaining safety constraints, achieving 86.4% terminal agreement with clinicians compared to IQL’s 75%. Our results suggest that geometric pessimism provides the necessary regularisation to safely overcome local optima in critical, real-world decision systems.
@@ 1 Introduction
Reinforcement Learning (RL) has achieved superhuman performance in simulated domains, solving complex strategy games such as Go and Dota 2 (Berner et al., 2019; Silver et al., 2017). In these environments, RL agents learn through an "online" paradigm, i.e. learning occurs through continuous interaction with the environment, where data is generated sequentially as a result of the agent’s own actions and is then immediately used to update the policy (Murphy, 2024). If an agent crashes a virtual car or loses a game then the simulation is simply reset. However, this reliance on trial-and-error has rendered RL effectively useless in high-stakes, real-world domains such as robotic surgery, sepsis management and autonomous driving (Levine et al., 2020). In these domains a single exploratory action no longer leads to a trivial "Game Over" screen, but instead may cause catastrophic hardware failure and even patient mortality. To enable safe deployment in such high-stakes settings, RL must fundamentally adapted beyond its trial-and-error paradigm.
To address this safety gap a new paradigm known as Offline RL has emerged. Instead of iteratively interacting with the world in a dynamic setting, Offline RL algorithms learn policies entirely from static, historical datasets (Levine et al., 2020). Such as previously collected hospital logs (e.g. MIMIC-III Critical Care dataset) or human driving data for the purpose of autonomous driving.
However, Offline RL introduces a subtle but dangerous failure mode, the problem of Distributional Shift (Levine et al., 2020), that results in value overestimatin which can have dire consequences. In classical RL, algorithms assume that optimistic estimates for untried actions are acceptable because the agent can directly attempt them by interacting with the environment and receive its respective reward. If the reward is low, the algorithm learns not to attempt that action again.
On the contrary, in a closed dataset setting, this learning process breaks down because the agent cannot interact with the environment to correct its mistakes during training. When the policy attempts an action outside its training distribution (out-of-distribution, OOD), it has the potential to hallucinate and assign arbitrarily high rewards simply because no data exists to suggest otherwise (Kumar et al., 2019). Since the agent cannot gather new information, these errors persist and propagate via the Bellman backup which is the foundational core of RL, leading to permanent overestimation of poor actions. Consequently, the agent assumes that unknown actions are the best actions (due to its erroneously high estimated value) causing it to confidently execute dangerous, erratic behaviours in the real world (Kumar et al., 2019).
@@ 2 Background
Conservative Q-Learning (CQL): Representing the Pessimistic camp, CQL is currently the gold standard for safety due to its conservative actions and is canonically a very strong baseline when researching new methods. It operates by adding a regularisation term to the objective function that actively "pushes down" the Q-Values of OOD actions (Kumar et al., 2020). Although effective at preventing overestimation, this comes at a very high computational cost as CQL requires sampling unseen actions during every training step to compute the log-sum-exp regulariser, resulting in slower convergence as well (Kumar et al., 2020).
Implicit Q-Learning (IQL): To address the computational inefficiency of CQL, IQL avoids querying OOD actions altogether (Kostrikov et al., 2021). Instead it treats the value function as a regression problem to estimate the Q value. It estimates the optimal Q-value functions via expectile regression, allowing IQL to remain entirely "in-sample" offering significant speed advantages and stability (Kostrikov et al., 2021). However, IQL lacks an explicit mechanism to detect OOD data. As a result in "fractured" dataset where gaps in the state space coverage exist, IQL could struggle to estimate the optimal Q-Values, leading to sub-par performance.
@@ 4 Methodology
We propose Geo-IQL (Geometric Implicit Q-Learning), an algorithm designed to inject geometric safety constraints into the already efficient IQL algorithm. The algorithm operates in a four-stage pipeline: (1) Embedding, (2) Geometric Uncertainty Estimation, (3) Robust Standardization, and (4) Reward Penalization.
@@ 4.1 State-Action Embedding
To measure how "familiar" a state-action pair is (i.e. taking a specific action in a given state), we first map the (state, action) data into a joint embedding space. We define the embedding vector $\phi(s,a)$ by concatenating the normalized state vector with the raw action vector:
@@
Design Decision: We explicitly excluded action normalization from this embedding. In control tasks, the magnitude of the action (e.g. how hard a motor spins) carries physical significance. Preserving the raw action scale ensures that Euclidean distance in this space reflects true physical divergence in control inputs.
Figure 1: 3-Dimensional visual of using geometry as a proxy for epistemic uncertainty.
@@ 4.3 Robust Standardisation
Raw Euclidean distances vary wildly depending on the environment dimensions. To make our algorithm transferable across different tasks (e.g. robotics vs. healthcare, as we later explore in this work), we must standardise these scores. Standard mean-variance normalisation is susceptible to outliers. Therefore, we engineer a robust standardisation utilising Median Absolute Deviation (MAD).
We define a "safe threshold", $\tau$ as the $\alpha$-quantile of distances in the dataset (set to $\alpha=0.3$ based on preliminary tuning). This explicitly categorizes the closest 30% of the data as the "Safe Core", receiving a penalty of 0. This then implies that the remaining 70% of the dataset is subject to penalisation. However the our standardisation ensures this is not a binary punishment, but a graded risk surface. We compute the robust spread $\sigma_{MAD}$ with $\epsilon$ to ensure division by zero does not occur:
@@
The standardized uncertainty score $U(s,a)$ is then:
@@
4.4 Density-Adaptive Penalization
@@
This factor decays from 1 to 0 as uncertainty increases. We use this to modulate the penalty weight $\lambda_{adapt}$:
@@
Figure 2: Visualizing the Adaptive Safety Mechanism. The blue cloud represents the training data manifold. Green Point: A query inside the dense region. The algorithm detects near-zero distance to neighbours and applies no penalty. Yellow Point: A query slightly off-manifold. This triggers the adaptive lambda, applying a mild correctional penalty. Red Point: A query far from the data. Effectively, OOD. The algorithm measures a large mean Euclidean distance from its neighbours and applies a heavy penalty to reject this action entirely.
@@ 4.6.1 IQL: Target and Loss
IQL utilises Expectile Regression to learn the value function without querying out-of-sample actions. The framework consists of Twin Q-Networks $Q_{\theta_{1,2}}$, a Value network $V_{\psi}$, and an Actor $\pi_{\phi}$. The standard critic objective minimizes the Mean Squared Error (MSE) against a target constructed from the value network:
$y=r(s,a)+\gamma V_{\Phi}(s^{\prime})$
@@ 4.6.2 Geo-IQL: Modifications
Crucially, the Value ($V_{\psi}$) and Actor ($\pi_{\phi}$) update rules remain identical to the original IQL formulation. This ensures that the geometric module acts as a strictly additive constraint on the value estimation, preserving the stability of the underlying algorithm. The entire algorithm is presented in Algorithm 1
@@
  Normalize: $U\leftarrow(\tilde{U}-\tau)/\sigma$
@@ 5.1 Robotic Locomotion (D4RL)
Environment: We evaluate on the D4RL MuJoCo suite (Fu et al., 2020), the standard benchmark for continuous control in offline RL. Specifically, we focus on Hopper, Walker2d, and HalfCheetah (see Fig. 3,4,5) (Park and Lee, 2024) which require controlling articulated robots to achieve stable locomotion. The tasks are as follows:
@@
Each of these robots are controlled by a policy recovered from the static dataset by the algorithm. Performance over these locomotion environments are reported as a Normalized Score, where 0 corresponds to a random policy and 100 to an expert policy. Scores exceeding 100 indicate the agent has outperformed the data-collecting policy that the dataset was generated from.
@@ 5.2 Sepsis Management (MIMIC-III)
Environment: Additionally to assess performance in a high-stakes, real-world setting we utilise the MIMIC-III Sepsis Cohort (Goldberger et al., 2000; Johnson et al., 2025, 2018, 2023). The objective is to prescribe optimal dosages of vasopressors and intravenous fluids (discrete action space $|\mathcal{A}|=25$) to maximise patient survival. Since online evaluation is impossible, as we do not test on live patients. We report metrics across three categories: Clinician Agreement, Policy Quality and Safety.
@@ 5.2.1 Clinician Agreement
    Probability of Clinician Action: At what probability would the policy take the action that the clinician took. Essentially representing how reasonable the policies believes that the clinician’s actions were. High probability signifies that the action were very reasonable and a low probability means that the action were highly suboptimal from the perspective of the policy
@@ 5.2.2 Policy Quality
    Terminal State Agreement: This measures how often the policy matched the clinician’s actions in the last few steps. In Sepsis treatment, the "middle" of a trajectory is often noisy as there are many ways to keep a stable patient stable. However, the end is critical. If the patient survived, then clinician’s final actions likely stabilised the patient enough for discharge. A high agreement here signifies that the policy learned "how to successfully discharge a patient". On the other hand, if the patient died, the clinician was likely fighting a crashing patient. High agreement here signifies that the policy successfully recognised a medical crisis and attempted actions that a clinician would deem the best.
@@ 5.3 Implementation and Hardware Constraints
All algorithms were implemented using the CORL (Clean Offline RL) library (Tarasov et al., 2023, 2022), to ensure reproducibility. Link to codebase can be found in the appendix. All experiments were conducted on a single consumer-grade laptop GPU, NVIDIA RTX 5060. This constraint demonstrates that Geo-IQL’s $\mathcal{O}(1)$ pre-computed mechanism makes safe AI accessible without requiring industrial-scale compute clusters (unlike ensemble methods). Additionally we report the mean and standard deviation over 5 random seeds to ensure statistical significance.
@@ 6.1 D4RL MuJoCo
As shown in Table 1, Geo-IQL achieves substantial gains in the fractured hopper-medium-replay-v2 domain, outperforming standard IQL by over 18 points while reducing inter-seed variance by approximately $\mathbf{4\times}$. Notably, Geo-IQL also surpasses the conservative baseline CQL, yet maintains the computational efficiency of IQL, avoiding the expensive iterative regularisation that is inherent to CQL.
@@
In the more stable multi-legged locomotion tasks (halfcheetah, walker2d), Geo-IQL achieves performance statistically equivalent to IQL, with overlapping confidence intervals. This behavior is expected and desirable. It confirms that our geometric penalty is adaptive, applying necessary conservatism in sparse, unstable regimes, such as Hopper’s hyper-sensitive survival dynamics where smallest errors and overestimation leads to early termination, all while not degrading performance in the denser, stable manifolds of HalfCheetah.
@@ 6.2 MIMIC-III Sepsis
In contrast, Geo-IQL demonstrates targeted optimisation. While it agrees less often on intermediate steps, reflected by its active adjustment of vasopressor dosages to maximize return. It achieves a positive value improvement, $\Delta Q=+0.0138$. Furthermore, from Fig 7. we see that Q-Improvement for Geo-IQL increases and tapers of (as its estimate of clinician slowly catches up to the policies estimate of itself) but is still positive. On the other hand, IQL is always negative.
Crucially, this optimization does not come at the cost of safety. Geo-IQL achieves a significantly higher Terminal State Agreement (86.39% vs 75.02%), indicating that while it may optimize the treatment path, it aligns far more robustly with clinician decisions during critical patient outcomes (discharge or crisis). The lower entropy of 0.69 further suggests that the geometric penalty helps the policy converge to a more confident and consistent treatment regime.
@@ 7 Conclusion
    Stability on Fractured Manifolds: In hyper sensitive dynamics of Hopper-Medium-replay-v2, where standard algorithms wildly oscillate between sustained success and instant failure, Geo-IQL stabilises learning, while outperforming computationally expensive baselines with the speed of IQL.
@@
    Efficiency: Unlike iterative regularisation methods like CQL that require sampling every training steps our precomputed lookup approach maintains a $\mathcal{O}(1)$ runtime complexity, making it scalable for large high-dimensional real-world datasets.

### v2

@@ Abstract
Offline Reinforcement Learning (RL) promises the recovery of optimal policies from static datasets, yet it remains susceptible to the overestimation of out-of-distribution (OOD) actions, particularly in fractured and sparse data manifolds. Current solutions necessitate a trade-off between computational efficiency and performance. Methods like CQL offer rigorous conservatism but require tremendous compute power while efficient expectile-based methods like IQL often fail to correct OOD errors on pathological datasets, collapsing to Behavioural Cloning. In this work, we propose Geometric Pessimism, a modular, compute-efficient framework that augments standard IQL with density-based penalty derived from $k$-nearest-neighbour distances in the state-action embedding space. By pre-computing the penalties applied to each state-action pair, our method injects OOD conservatism via reward shaping with a $\mathcal{O}(1)$ training overhead to the training loop.
[…]
@@ 1 Introduction
Reinforcement Learning (RL) has achieved superhuman performance in simulated domains, solving complex strategy games such as Go and Dota 2 (Berner et al., 2019; Silver et al., 2017). In these environments, RL agents learn through an “online” paradigm, i.e. learning occurs through continuous interaction with the environment, where data is generated sequentially as a result of the agent’s own actions and is then used to update the policy (Murphy, 2024). If an agent crashes a virtual car or loses a game then the simulation is simply reset. However, this reliance on trial-and-error has rendered RL effectively useless in high-stakes, real-world domains such as robotic surgery, sepsis management and autonomous driving (Levine et al., 2020). In these domains a single exploratory action no longer leads to a trivial “Game Over” screen, but instead may cause catastrophic hardware failure and even patient mortality. To enable safe deployment in such high-stakes settings, RL must fundamentally adapt beyond its trial-and-error paradigm.
To address this safety gap a new paradigm known as Offline RL has emerged. Instead of iteratively interacting with the world in a dynamic setting, Offline RL algorithms learn policies entirely from static, historical datasets (Levine et al., 2020), such as previously collected hospital logs (e.g. MIMIC-III Critical Care dataset) or human driving data for the purpose of autonomous driving.
However, Offline RL introduces a subtle but dangerous failure mode, the problem of Distributional Shift (Levine et al., 2020), that results in value overestimation which can have dire consequences. In classical RL, algorithms assume that optimistic estimates for untried actions are acceptable because the agent can directly attempt them by interacting with the environment and receive its respective reward. If the reward is low, the algorithm learns to reduce the probability of attempting that action.
On the contrary, in a closed dataset setting, this learning process breaks down because the agent cannot interact with the environment to correct its mistakes during training. When the policy attempts an action outside its training distribution (out-of-distribution, OOD), it has the potential to hallucinate and assign arbitrarily high rewards simply because no data exists to suggest otherwise (Kumar et al., 2019). Since the agent cannot gather new information, these errors persist and propagate via the Bellman backup, which is the foundational core of RL, leading to permanent overestimation of poor actions. Consequently, the agent assumes that unknown actions are the best actions (due to its erroneously high estimated value) causing it to confidently execute dangerous, erratic behaviours in the real world upon deployment (Kumar et al., 2019).
@@ 2 Background
Conservative Q-Learning (CQL): Representing the Pessimistic camp, CQL is currently the gold standard for safety due to its conservative actions and is canonically a very strong baseline when researching new methods. It operates by adding a regularisation term to the objective function that actively “pushes down” the Q-Values of OOD actions (Kumar et al., 2020). Although effective at preventing overestimation, this comes at a very high computational cost as CQL requires sampling unseen actions during every training step to compute the log-sum-exp regulariser, resulting in slower convergence as well (Kumar et al., 2020).
Implicit Q-Learning (IQL): To address the computational inefficiency of CQL, IQL avoids querying OOD actions altogether (Kostrikov et al., 2021). Instead it treats the value function as a regression problem to estimate the Q value. It estimates the optimal Q-value functions via expectile regression, allowing IQL to remain entirely “in-sample” offering significant speed advantages and stability (Kostrikov et al., 2021). However, IQL lacks an explicit mechanism to detect OOD data. As a result in “fractured” dataset where gaps in the state space coverage exist, IQL could struggle to estimate the optimal Q-Values, leading to sub-par performance.
@@ 4 Methodology
We propose Geo-IQL (Geometric Implicit Q-Learning), an algorithm designed to inject geometric safety constraints into the already efficient IQL algorithm. The algorithm operates in a four-stage pipeline: (1) Embedding, (2) Geometric Uncertainty Estimation, (3) Robust Standardisation, and (4) Reward Penalisation.
@@ 4.1 State-Action Embedding
To measure how “familiar” a state-action pair is (i.e. taking a specific action in a given state), we first map the (state, action) data into a joint embedding space. We define the embedding vector $\phi(s,a)$ by concatenating the Gaussian-normalised state vector with the raw action vector:
@@
Design Decision: We explicitly excluded action normalisation from this embedding. In control tasks, the magnitude of the action (e.g. how hard a motor spins) carries physical significance. Preserving the raw action scale ensures that Euclidean distance in this space reflects true physical divergence in control inputs.
Figure 1: 3-Dimensional visualisation of using geometry as a proxy for epistemic uncertainty.
@@ 4.3 Robust Standardisation
Raw Euclidean distances vary wildly depending on the environment dimensions. To make our algorithm transferable across different tasks, we must standardise these scores. Standard mean-variance normalisation is susceptible to outliers. Therefore, we engineer a robust standardisation utilising Median Absolute Deviation (MAD).
We define a “safe threshold”, $\tau$, as the $\alpha$-quantile of distances in the dataset ($\alpha=0.3$ based on preliminary tuning). This explicitly categorises the closest 30% of the data as the “Safe Core”, receiving a penalty of 0. This then implies that the remaining 70% of the dataset is subject to penalisation. However the our standardisation ensures this is not a binary punishment, but a graded risk surface. We compute the robust spread $\sigma_{MAD}$ with a small $\epsilon$ to ensure division by zero does not occur:
@@
The standardised uncertainty score $U(s,a)$ is then:
@@
4.4 Density-Adaptive Penalisation
@@
This factor decays from 1 to 0 as uncertainty increases. We use this to modulate the penalty weight $\lambda_{adapt}$ from $\lambda_{base}$ a hyper-parameter:
@@
Figure 2: Visualising the Adaptive Safety Mechanism. The blue cloud represents the training data manifold. Green Point: A query inside the dense region. The algorithm detects near-zero distance to neighbours and applies no penalty. Yellow Point: A query slightly off the manifold. This triggers the adaptive lambda, applying a mild correctional penalty. Red Point: A query far from the data. Effectively, OOD. The algorithm measures a large mean Euclidean distance from its neighbours and applies a heavy penalty to reject this action entirely.
@@ 4.6.1 IQL: Target and Loss
IQL utilises Expectile Regression to learn the value function without querying out-of-sample actions. The framework consists of Twin Q-Networks $Q_{\theta_{1,2}}$, a Value network $V_{\psi}$, and an Actor $\pi_{\phi}$. The standard critic objective minimises the Mean Squared Error (MSE) against a target constructed from the value network:
$y=r(s,a)+\gamma V_{\psi}(s^{\prime})$
@@ 4.6.2 Geo-IQL: Modifications
Crucially, the Value ($V_{\psi}$) and Actor ($\pi_{\phi}$) update rules remain identical to the original IQL formulation. This ensures that the geometric module acts as a strictly additive constraint on the value estimation, preserving the stability of the underlying algorithm. The entire algorithm is presented in Algorithm 1.
@@
  Normalise: $U\leftarrow(\tilde{U}-\tau)/\sigma$
@@ 5.1 Robotic Locomotion (D4RL)
Environment: We evaluate on the D4RL MuJoCo suite (Fu et al., 2020), the standard benchmark for continuous control in offline RL. Specifically, we focus on Hopper, Walker2d, and HalfCheetah (see Figures 3, 4 and 5) (Park and Lee, 2024) which require controlling articulated robots to achieve stable locomotion. The tasks are as follows:
@@
Each of these robots are controlled by a policy recovered from the static dataset by the algorithm. Performance over these locomotion environments are reported as a Normalised Score, where 0 corresponds to a random policy and 100 to an expert policy. Scores exceeding 100 indicate the agent has outperformed the data-collecting policy that the dataset was generated from.
@@ 5.2 Sepsis Management (MIMIC-III)
Environment: Additionally, to assess performance in a high-stakes, real-world setting we utilise the MIMIC-III Sepsis Cohort (Goldberger et al., 2000; Johnson et al., 2025, 2018, 2023). The objective is to prescribe optimal dosages of vasopressors and intravenous fluids (discrete action space $|\mathcal{A}|=25$) to maximise patient survival. Since online evaluation is impossible, as we do not test on live patients. We report metrics across three categories: Clinician Agreement, Policy Quality and Safety.
@@ 5.2.1 Clinician Agreement
    Probability of Clinician Action: At what probability would the policy take the action that the clinician took? This essentially represents the policy’s belief on how reasonable the clinician’s actions were. High probability signifies that the action were very reasonable and a low probability means that the action were highly suboptimal from the perspective of the policy
@@ 5.2.2 Policy Quality
    Terminal State Agreement: This measures how often the policy matched the clinician’s actions in the last few steps. In Sepsis treatment, the “middle” of a trajectory is often noisy as there are many ways to keep a stable patient stable. However, the end is critical. If the patient survived, then clinician’s final actions likely stabilised the patient enough for discharge. A high agreement here signifies that the policy learned “how to successfully discharge a patient”.
    On the other hand, if the patient died, the clinician was likely responding to acute clinical deterioration. High agreement here signifies that the policy successfully recognised a medical crisis and attempted actions that a clinician would deem the best.
@@ 5.3 Implementation and Hardware Constraints
All algorithms were implemented using the CORL (Clean Offline RL) library (Tarasov et al., 2023, 2022), to ensure reproducibility. Link to codebase can be found in the appendix. All experiments were conducted on a single consumer-grade laptop GPU, NVIDIA RTX 5060. This constraint demonstrates that Geo-IQL’s $\mathcal{O}(1)$ access to pre-computed penalty mechanism makes safe AI accessible without requiring industrial-scale compute clusters (unlike ensemble methods). Additionally we report the mean and standard deviation over 5 random seeds to ensure statistical significance.
@@ 6.1 D4RL MuJoCo
As shown in Table 1, Geo-IQL achieves substantial gains in the fractured hopper-medium-replay-v2 domain, outperforming standard IQL by over 18 points while reducing inter-seed standard deviation by approximately $\mathbf{4\times}$. Notably, Geo-IQL also surpasses the conservative baseline CQL, yet maintains the computational efficiency of IQL, avoiding the expensive iterative regularisation that is inherent to CQL.
@@
In the more stable multi-legged locomotion tasks (halfcheetah, walker2d), Geo-IQL achieves performance statistically equivalent to IQL, with overlapping standard deviation. This behavior is expected and desirable. It confirms that our geometric penalty is adaptive, applying necessary conservatism in sparse, unstable regimes, such as Hopper’s hyper-sensitive survival dynamics where the smallest errors and overestimation leads to early termination, all while not degrading performance in the denser, stable manifolds of HalfCheetah.
@@ 6.2 MIMIC-III Sepsis
In contrast, Geo-IQL demonstrates targeted optimisation. While it agrees less often on intermediate steps, reflected by its active adjustment of vasopressor dosages to maximise return. It achieves a positive value improvement, $\Delta Q=+0.0138$. Furthermore, from Fig 7. we see that Q-Improvement for Geo-IQL increases and tapers of (as its estimate of clinician slowly catches up to the policy’s estimate of itself) but is still positive. On the other hand, IQL is always negative.
Crucially, this optimisation does not come at the cost of safety. Geo-IQL achieves a significantly higher Terminal State Agreement (86.39% vs 75.02%), indicating that while it may optimise the treatment path, it aligns far more robustly with clinician decisions during critical patient outcomes (discharge or crisis). The lower entropy of 0.69 further suggests that the geometric penalty helps the policy converge to a more confident and consistent treatment regime.
@@ 7 Conclusion
    Stability on Fractured Manifolds: In hyper-sensitive dynamics of Hopper-Medium-replay-v2, where standard algorithms wildly oscillate between sustained success and instant failure, Geo-IQL stabilises learning, while outperforming computationally expensive baselines with the speed of IQL.
@@
    Efficiency: Unlike iterative regularisation methods like CQL that require sampling every training steps our precomputed lookup approach maintains a $\mathcal{O}(1)$ runtime overhead, making it scalable for large high-dimensional real-world datasets.